OpenStudy (thecatman):

The owner of a large company is conducting a survey about job satisfaction, including questions about salary, hours, stress, and other conditions. There are about eight hundred employees in the company, including one hundred executive positions, six hundred fifty middle-management positions, and fifty custodial positions. The owner wants to include about ten percent of his employees in the survey sample. Identify a sampling method that would lead to a representative sample for the survey. Explain why it would be a good choice and give details about the process.

OpenStudy (amistre64):

reposting the same question over and over again is considered a form of spam ... just in case you were not aware if that

OpenStudy (thecatman):

it is just hard

OpenStudy (amistre64):

which sampling defintion splits a population into subgroup, then takes samples from the subgroups?

OpenStudy (amistre64):

if you add up all the positions, to determine how much is 10% ... you would have a basis for the subgroups

OpenStudy (thecatman):

10% of 100 = 10 executives 10% of 650 = 65 middle managers 10% of 50 = 5 custodial workers

OpenStudy (thecatman):

is this right

OpenStudy (amistre64):

oh, they way there are 800 positions; 10% is 80 100/800 = 10/80 pick 10 from that group 650/800 = 65/80 pick 65 loks good

OpenStudy (amistre64):

the name of that sampling is called what tho?

OpenStudy (thecatman):

?

OpenStudy (amistre64):

when we divide the population into different groups; its called stratification. then picking a simple random sample from each group gives us a good representative sample the one that often gets confused with it is the cluster; the cluster separates and then takes one of the clusters as the sample

OpenStudy (amistre64):

good luck with it, i have to run